A hot air balloon is rising vertically at 9m/s. At an altitude of 500m, a ballast bag with a mass of 25kg breaks loose and falls under the influence of gravity. How long will it take to reach the ground and how fast will it be going when it hits the ground assuming no air resistance. Use g = 10 m/s2. Be sure to explain your reasoning while using our standard 4-point problem solving approach.
Answer #1

In vertical:

mass was with ballon so it would be having initial velocity same as ballon.

vi = 9 m/s

a = - 10 m/s^2

r = - 500 m   (as it have to go down to reach ground )

using r = vi*t + at^2 /2

-500 = 9t - 10t^2/2

5t^2 - 9t - 500 = 0

t = 10.94 s .....Ans


vf = vi + at

vf = 9 - 10*10.94

vf = - 100.4 m/s